A Telkom company bond carries an 8% coupon, paid semi-annually. The par



value is R1000, and the bond matures in six years. If the bond currently sells for



R911.37, what is its yield to maturity?

1
Expert's answer
2022-03-14T17:25:38-0400

Solution


Number of periods "=\frac{2 \ periods }{year}\times (number\ of \ years)"


"=\frac{2 \ periods }{year}\times (6 \ years)= 12"


Par/ Face Value "=R1000"


Since the company bond carries an "8\%" coupon, therefore,


"PMT=FV\times \frac{8\%}{2}=R40"


Current selling Price of Bond "= R 911.37"


Therefore, using


"911.37 = 40 \times \frac{{\left[ {1 - \frac{1}{{{{\left( {1 + r} \right)}^{12}}}}} \right]}}{r} + \frac{{1000}}{{{{\left( {1 + r} \right)}^{12}}}}"


Solving for yield rate "r", we get


"r=0.1=10\%"


Therefore, yield to maturity is "10\%"